OpenStudy (anonymous):

raul invests $1000 in a savings account that earns 5% interest compounded annually. how much money will he have in the account at the end of 5 years

OpenStudy (anonymous):

1000*(1.05)^5

OpenStudy (anonymous):

how often does the account earn intrest? monthly or yearly or what?

OpenStudy (anonymous):

Since, it is compounded anually n ( R ) Amount = P + (1 + ---- ) ( 100) where P = sum deposited -------- $1000 R = rate of interest -------- 5% n = time in years -------- 5 years Amount = 1000 (1 + 0.05)^5 = 1000 (1.05)^5 = 1000 * 1.276 = 1276 so at end of 5 years, he will have $ 1276 in his account.

OpenStudy (anonymous):

thanks Angela :D

OpenStudy (anonymous):

where do u live harkirat?

OpenStudy (keenan):

i got 4000

OpenStudy (anonymous):

I live in capital of India, Delhi..... and U???

OpenStudy (angela210793):

UW (U taught me how to solve this kind of problem) Thnx :)

OpenStudy (anonymous):

I am glad that u hv LEARNT it now!!! Many just keep coming back with same problems again and again... That is why i give answer in detail so that the asker can learn to do it himself/herself.......

OpenStudy (anonymous):

@him1618 ???????

OpenStudy (angela210793):

well i sometimes stop answering to ppl who ask same kind of q 10000000000000000000000000000000 times :P:P ur patient ...good 4 u :P :)

OpenStudy (angela210793):

@Harkirat take a look at this pls http://openstudy.com/users/angela210793#/users/angela210793/updates/4e058ed40b8bd74af49b6b8f

OpenStudy (anonymous):

u talking abt the video????

OpenStudy (angela210793):

aha...

OpenStudy (anonymous):

yes, it is nice but 10 Miniutes long !!!!

OpenStudy (angela210793):

lol :P yea I loved crossing over part :P:P

OpenStudy (anonymous):

yes it is really informative as well as entertaining......... :D

OpenStudy (angela210793):

^_^ :)